OpenStudy (anonymous):

A line on a coordinate grid passes through the points (-1, -3) and (4, 5). What is the slope of a line perpendicular to this line? Write the answer as a fraction in lowest terms.

OpenStudy (danjs):

can you find the slope of the line through those 2 points?

OpenStudy (anonymous):

how ?

OpenStudy (danjs):

recall: slope = change in y value / change in x value = (y2-y1) / x2-x1)

OpenStudy (danjs):

Where a points given are (x1,y1) and (x2,y2)

OpenStudy (danjs):

\[slope = \frac{ 5 -(-3) }{ 4-(-1) } = ?\]

OpenStudy (danjs):

what is the slope of the given line with those 2 points?

OpenStudy (anonymous):

8/5 ?

OpenStudy (danjs):

ok correct, so we need a line that is perpendicular to that slope, to find the slope of that line, flip the slope over and make it negative (negative recipral) What is the slope of the perpendicular line?

OpenStudy (anonymous):

-8/5?

OpenStudy (danjs):

close: you have to flip the fraction upside down, AND take the negative

OpenStudy (danjs):

THen, use that slope of the perpendicular line m = (-5/8) and one of the points in the questin,(x1,y1), to form a equation of the perpendicular line, y - y1 = m(x - x1)